Consider the given circuit. Given I2 = 2 A and R = 35 Ω, determine the value of Is.Consider the given circuit. Given 12 = 2 A and R = 351, determine the value of Is. llll llll $ 125 22 (1 The value of Is =
Answer #1

Answer - KAZI I qui 10 As In L = - 13 I= -5 Iz Ij = -5x 2 I=-10 Amp Va = - RT, -35 (-10) And Vo = 350 Volt 2AS V2 = -5 V₂ = -5 ui V₂ = -5* 350 V2 = -1750 volt Applying nodal analysis at node 0- V = Ist In -1750 T25 . Is + I2 Is = -14